### v2

# Sequential Model for Predicting Patient Adherence in Subcutaneous Immunotherapy for Allergic Rhinitis

Abstract
Objective: Subcutaneous Immunotherapy (SCIT) is the long-lasting causal treatment of allergic rhinitis. How to enhance the adherence of patients to maximize the benefit of allergen immunotherapy (AIT) plays a crucial role in the management of AIT. This study aims to leverage novel machine learning models to precisely predict the risk of non-adherence of patients and related systematic symptom scores, to provide a novel approach in the management of long-term AIT.
Methods: The research develops and analyzes two models, Sequential Latent Actor-Critic (SLAC) and Long Short-Term Memory (LSTM), evaluating them based on scoring and adherence prediction capabilities.
Results: Excluding the biased samples at the first time step, the predictive adherence accuracy of the SLAC models is from $60\,\%$ to $72\%$, and for LSTM models, it is $66\,\%$ to $84\,\%$, varying according to the time steps. The range of Root Mean Square Error (RMSE) for SLAC models is between $0.93$ and $2.22$, while for LSTM models it is between $1.09$ and $1.77$. Notably, these RMSEs are significantly lower than the random prediction error of $4.55$.
Conclusion: We creatively apply sequential models in the long-term management of SCIT with promising accuracy in the prediction of SCIT nonadherence in Allergic Rhinitis (AR) patients. While LSTM outperforms SLAC in adherence prediction, SLAC excels in score prediction for patients undergoing SCIT for AR. The state-action-based SLAC adds flexibility, presenting a novel and effective approach for managing long-term AIT.
[…]
1 Keywords: Allergic rhinitis, Allergen immunotherapy, Adherence, Sequential model, Latent variable model
2 Introduction
Allergic rhinitis is characterized by allergen-specific IgE-mediated inflammation in upper respiratory inflammation with a prevalence of up to $30\,\%$ worldwide (Meltzer, 2016). In addition to allergen avoidance as the superior criterion, allergen-specific immunotherapy (AIT) aims to induce specific allergen immune tolerance, consequently achieving a status of clinical symptom remission. The repeatable intake of the specific unmodified or chemically modified allergens (allergoids) was the key to maintaining the symptoms. Among these approaches of AIT, subcutaneous immunotherapy (SCIT), sublingual immunotherapy (SLIT), and lymphatic immunotherapy (SLIT) are demonstrated as the mainstream treatments regarding efficacy, safety, and side effects. Compared to the SLIT, SCIT is a clinic-dependent treatment in which the patient accepted an allergen extract injection subcutaneously. It is divided into the initial treatment stage (dose accumulation stage) and the maintenance treatment stage (dose maintenance stage). The World Health Organization (WHO) recommends that immunotherapy be maintained for three to five years and clinically recommended for at least two years. Patient adherence is a critical factor in ensuring long-lasting efficacy and sustaining symptom-relieving effects.
Due to the long duration of SCIT, cumbersome process, slow onset, individual differences in treatment effect, and other factors fundamentally impact the completeness of therapeutics. From the reported studies on AIT, the rate of adherence ranged from around $25\%$ to over $90\%$ (Passalacqua et al., 2013). The WHO adopted the definition of ‘adherence’ as “the extent to which a person’s behavior, such as taking medication, following a diet, or executing lifestyle changes, corresponds with agreed recommendations from a health care provider” (Organization et al., 2003). In recent EAACI guidelines, it is highlighted to educate patients on how immunotherapy works and on explaining the importance of compliance to the regular doses for three years of treatment(Roberts et al., 2018).
The multiple approaches were introduced into the field of improving adherence and supervising patient outcomes with systematic and technological interventions to prevent incomplete discontinuation of the treatment. The intervention from a clinic in advance running through the whole treatment cycle was approved as an effective approach. In facing the multitude of personalized data from patients, how to precisely identify and assess the risk of upcoming non-adherent behavior, a clinical prediction model is promising in the application.
In healthcare, machine learning, especially sequential models, stands at the forefront of innovation, providing new ways to analyze complex medical data and improve patient treatments. These models excel in processing and analyzing time-dependent data, making them ideal for predicting patient adherence to treatments like SCIT for AIT. By effectively utilizing sequential data, these algorithms uncover hidden temporal patterns and correlations, leading to more accurate and personalized treatment plans. In this study, we have selected and evaluated two specific sequential models tailored to this scenario. Our findings demonstrate that these models are not only effective in predicting patient adherence to medical treatments but also invaluable in enhancing treatment strategies, thereby making a significant contribution to patient-centered healthcare.
3 related work
Non-sequential models for adherence prediction. A hybrid model was presented to combine artificial neural networks (ANNs) and the genetic algorithm (GA) to examine 26 factors influencing adherence to dietary plans (Mousavi et al., 2022). Wang et al. (2020) developed two machine learning methods, neural network (NN) and support vector machine (SVM), to predict nonadherence in Crohn’s Disease patients. These models assist caregivers by streamlining the intervention process using azathioprine (AZA). Warren et al. (2022) investigated the relationship between patient adherence to prescribed medication for opioid use disorder and various risk factors in patients with opioid use disorder. Four different machine learning classifiers were proposed, including Logistic Regression, Decision Tree, Random Forest, and XGBoost. They compared the performance of these classifiers to explore the association. Since the adherence threshold can determine the prediction model’s performance by the binary adherence status, Ruff et al. (2019) proposed a simulation framework to assess the performance of prediction models based on different adherence thresholds.
Sequential models for adherence prediction. Hsu et al. (2022) investigated the advantages of incorporating patient history into the prediction of medication adherence. They assessed the performance of temporal deep learning models, particularly LSTM and simple Recurrent Neural Networks (RNN), and compared these with non-temporal models such as Multilayer Perceptrons (MLP), ridge classifiers, and logistic regression. To optimize the efficacy of cognitive training for older adults, Singh et al. (2022) employed multivariate time series analysis and developed personalized models for each patient to capture their unique adherence patterns. However, the sequential data of patients is often characterized by fluctuating adherence and high dropout rates, resulting in uneven, unaligned, and missing values in the time series data. To address this challenge, Schleicher et al. (2023) applied change point detection to identify phases with varying dropout rates, presented methods for handling uneven and misaligned time series, and utilized time series classification to predict the user’s phase.
4 Methods
In this section, we outline the comprehensive methodologies employed in our study. These methods are designed to rigorously analyze the effectiveness of SCIT in treating AR patients. We detail the study design, population criteria, treatment protocols, and sequential models to provide a clear understanding of our research framework.
4.1 Study design
The study design is a critical component that shapes the direction and reliability of our research. It includes a systematic approach to selecting the study population, the treatment methods applied, and the evaluation criteria.
4.1.1 Population
[…]
4.1.2 SCIT treatment and evaluation
[…]
Patients receive regular treatment evaluations, including symptom scores and medication scores. The symptom score recorded a total of nasal symptoms (nasal itching, sneezing, rhinorrhea, nasal congestion), ocular symptoms (ocular itching, lacrimation), and pulmonary symptoms (shortness of breath, tightness in chest, perennial cough, wheezing), and assessed symptom severity using the visual analogue scale (VAS). In the VAS score, the score of each symptom is from 0 to 10. 0 indicates that the patient has no discomfort and 10 indicates that the patient is very uncomfortable. The patient gives the score of each symptom according to the actual situation, and the sum of all symptom scores is the symptom score. Medication score recorded the use of current adjuvant medication within 1 month to reach symptom relief. The use of oral antihistamines, antileukotrienes, and bronchodilators were recorded as one point, local glucocorticoids as two points, oral glucocorticoids or combined medication (hormones and $\beta 2$ receptor agonists) as three points, and the total cumulative score was the medication score. Symptom scores and medication scores were assessed once at registration of SCIT and then thereafter.
Due to the separated injection regimen within 16 weeks and thereafter, all the chosen patients completed the four months of SCIT, we chose the fourth month as the starting point of the observation. According to our previous experience, one year after the start was the peak of the withdrawal, so we added a time point at 18 months to further assess and follow up on the related symptom score and individual status. The data collection spans six time steps: at 0, 4, 12, 18, 24, and 36 months. This approach is standard in medical treatment, although for optimal model performance, an equal distribution of time intervals would be preferable.
4.1.3 Data Collection
Data were collected from patient records in hospitals, and the following information was extracted for analysis: patient age, gender, distance to clinic, ratio of AIT cost to family income, allergen test results, etc., as well as patient VAS system score and medication score information, including baseline data of patients before injection therapy, adverse reactions to SCIT. For the descriptive analysis, categorical variables were given as numbers and percentages, and continuous variables were presented using mean, standard deviation, median, interquartile range (IQR), and minimum and maximum values.
4.1.4 Survey methods
Adherence was defined as the accomplishment of three years of AIT including the patients further received AIT. Non-adherence was defined as discontinuation of AIT at random time points during three years. The follow-up contents included (1) the main reasons for patients’ discontinuation of treatment; (2) the duration of discontinuation of treatment, and (3) Allergic symptoms after discontinuation of treatment.
4.2 Sequential models
The focus of our study is the development of sequential models that can efficiently and accurately predict the progression of symptoms and adherence in patients undergoing SCIT. This involves a comprehensive analysis of the data collected, structured to provide insights into the treatment’s effectiveness and patient compliance over time. Additionally, we explore and compare two distinct sequential models.
4.2.1 Data
variables patients
total adherent non-adherent
age $\leq 12$ 96 (46.7) 40 56
13–17 30 (14.6) 10 20
$\geq 18$ 79 (38.7) 23 56
gender Female 62 (30) 22 40
Male 143 (70) 51 92
distance to clinic(km) $\leq 10$ 136 (66) 56 80
$>10$ 69 (34) 17 52
cost/family income(%) $<30$ 107 (52.4) 37 70
30–50 77 (37.4) 32 45
$>50$ 21 (10.2) 4 17
[…]
EOS % 0.05; 0.04 0.05; 0.05 0.05; 0.05
$\Delta$NR(%) 16.67; 59.70 30.00; 92.80 14.80; 50.00
$\Delta$PNIF(%) 11.90; 34.50 12.70; 39.30 11.10; 28.80
total IgE (kU/L) 286; 543 340; 487 226; 555
sIgE of Derp (kU/L) 30.80; 68.480 31.30; 74.40 30.40; 67.80
sIgE of Derf (kU/L) 40.00; 68.20 40.60; 75.10 37.10; 65.70
Derp SPT SI 1.04; 0.58 1.00; 0.59 0.82; 0.55
Derf SPT SI 1.00; 0.50 0.82; 0.51 0.80; 0.45
Table 1: Demographic and clinical data of the patients under subcutaneous immunotherapy. In the rows from Age to Cost/Family income, values indicate the number of patients (percentage, if available). Other rows represent the median and IQR. P-values are omitted due to their large values.
reasons for SCIT withdrawal number of non-adherent patients
5–12 mths 13–18 mths 19–24 mths 25–36 mths total by reason
no clinical improvement 18 11 8 21 58
medical issue 3 1 2 0 6
improved efficacy 0 0 0 24 24
schooling 3 3 0 5 11
side effects 2 1 1 2 6
Covid-19 9 7 3 1 20
personal issue 0 3 0 4 7
total by time period 35 26 14 57 132
Table 2: Detailed reasons for withdrawal from SCIT at different time points.
Figure 1: Histogram of scores across six-time steps. Score value (horizontal axis) vs. count (vertical axis).
We have a dataset $D$, comprising sequences $x_{1},\dots,x_{T}\in R^{11}$, $y_{1},\dots,y_{T-1}\in R^{1}$, and a corresponding action $a_{t}\in R^{1}$. In the context of healthcare, the observations encompass $y_{t}$ (see Table 2) whether the patient will cease the treatment in the interval between the scoring measurements at $x_{t}$ and $x_{t+1}$. The actions $a_{t}$ represent the ongoing medical procedures for the patient during the period from $x_{t}$ to $x_{t+1}$ (see Fig. 1). In this context, $a_{t}$ is binary, reflecting whether treatment is given, and is numerically equivalent to the adherence variable $y_{t}$. Despite their numerical equivalence, we maintain a distinction between action $a_{t}$ and adherence $y_{t}$ to enhance model clarity and accommodate future research expansions, potentially allowing for a wider range of action values. For each patient, we possess basic information $s\in R^{14}$ which includes age, gender, commute distance to clinic, ratio of cost to family income, eosinophils count, eosinophils percentage, nasal allergen provocation test (change of nasal resistance, $\Delta\text{NR}(\%)$), peak nasal inspiratory flow. $\Delta\mathrm{PNIF}(\%)$), serum total IgE level, sIgE of Dermatophagoides pteronyssinus (Derp), sIgE of Dermatophagoides farinae (Derf), skin prick test (Derp, Derf) (see Table 1 for more details).
4.2.2 Sequential latent variable models
[…]
figures/lvm
Figure 2: SLAC model schematic. Solid and dashed lines denote the generative and inference model pathways, respectively. The gray circles represent observed data, and the white circles denote latent variables. The figure is adapted from (Lee et al., 2020).
In our research, we employ the Stochastic Latent Actor-Critic (SLAC) model (Lee et al., 2020). However, our application differs from the original use of SLAC which is typically associated with reinforcement learning. Instead, we leverage its sequential latent-variable model without Actor-Critic. This approach aligns with similar methodologies found in other works (Krishnan et al., 2015; Karl et al., 2016; Gregor et al., 2018). The choice of the SLAC model was motivated by its ability to facilitate more efficient learning and superior generalization in intricate environments.
The sequential latent variable model consists of an inference model and a generative model (see Fig. 2). The inference model in a sequential latent variable model typically aims to approximate the posterior distribution of the latent variables given the observed data. It tries to infer the hidden states $z$ based on the observed inputs $x$ and some internal state $s$. The inference models the probability distributions of the latent variables $z^{1}$ and $z^{2}$ at different time steps. $q_{\phi}$ denotes the variational distribution parameterized by $\phi$,
$z^{1}_{1}$ $\sim q_{\phi}(z_{1}^{1}|x_{1},s)$ (1)
$z^{2}_{1}$ $\sim p_{\phi}(z^{2}_{1}|z_{1}^{1})$ (2)
$z^{1}_{t+1}$ $\sim q_{\phi}(z^{1}_{t+1}|x_{t+1},z^{2}_{t},a_{t})$ (3)
$z^{2}_{t+1}$ $\sim p_{\phi}(z^{2}_{t+1}|z^{1}_{t+1},z^{2}_{t},a_{t}).$ (4)
The generative model, on the other hand, describes how the observed data is generated from the latent variables. It defines a process by which the latent states $z$ give rise to the observations $x$. The generative model is described by a set of equations that represent the probability distributions of both the initial latent states and their transitions over time, as well as the likelihood of the observations given the latent states, with $p_{\phi}$ indicating the parameterized generative distribution.
$z^{1}_{1}$ $\sim p(z_{1}^{1})$ (5)
$z^{2}_{1}$ $\sim p_{\phi}(z^{2}_{1}|z_{1}^{1})$ (6)
$z^{1}_{t+1}$ $\sim p_{\phi}(z^{1}_{t+1}|z^{2}_{t},a_{t})$ (7)
$z^{2}_{t+1}$ $\sim p_{\phi}(z^{2}_{t+1}|z^{1}_{t+1},z^{2}_{t},a_{t})$ (8)
$x_{t}$ $\sim p_{\phi}(x_{t}|z_{t}^{1},z^{2}_{t})$ (9)
$y_{t}$ $\sim p_{\phi}(y_{t}|z_{t}^{1},z^{2}_{t}).$ (10)
We have the evidence lower bound (ELBO):
$\log p_{\phi}(x_{1:t+1}|a_{1:t})\geq$ $\bigg{[}\mathop{\mathbb{E}}_{(x_{1:T},a_{1:T-1})\sim D}\bigg{[}\mathop{\mathbb{E}}_{z_{1:T}\sim q_{\phi}}\sum^{T-1}_{t=0}\Big{(}\log p_{\phi}(x_{t+1}|z_{t+1})$ (11)
$-D_{KL}(q_{\phi}(z_{t+1}|x_{t+1},z_{t},a_{t})\|p_{\phi}(z_{t+1}|z_{t},a_{t}))\Big{)}\bigg{]}\bigg{]}.$
For ease of notation, we have $q(z_{1}|x_{1},z_{0},a_{0}):=q(z_{1}|x_{1},s)$ and $p(z_{1}|z_{0},a_{0}):=p(z_{1})$. The ELBO provides a lower bound to the log-likelihood of the observed data, which is computationally intractable to compute directly. It is composed of two terms: the expected log-likelihood of the observed data given the latent variables, and the Kullback-Leibler (KL) divergence between the variational distribution and the prior distribution of the latent variables. Minimizing the KL divergence can be interpreted as enforcing the variational distribution to be as close as possible to the prior, while maximizing the expected log-likelihood ensures that the model accurately captures the distribution of the observed data. To predict the adherence, we have $\log p_{\phi}(y_{t+1}|z_{t+1})$ as a regularisor in the loss function.
The objective is to compute the parameters $\phi$ that minimize the KL divergence between the variational and prior distributions of the latent variables, subject to certain constraints. These constraints are related to the expected log-likelihood of the data under the model and are represented by the inequalities with thresholds $\xi$. These thresholds ensure that while minimizing the losses, the model also satisfies a minimum standard for score prediction and adherence classification performances.
Latent variable models, such as Variational Autoencoders (VAEs) (Kingma and Welling, 2014; Rezende et al., 2014) and their variants (e.g., SLAC), often encounter challenges (Sønderby et al., 2016; Kingma et al., 2016). Furthermore, a higher ELBO does not always lead to enhanced predictive performance, as discussed by Alemi et al. (2018); Higgins et al. (2017). However, the integration of scheduling strategies inspired by constrained optimization methods has been shown to significantly improve the training of latent variable models (Rezende and Viola, 2018; Klushyn et al., 2019). Consequently, we formulate the training of our model into an optimization problem
$\min_{\phi}$ $\mathop{\mathbb{E}}_{(x_{1:T},a_{1:T-1})\sim D}\left[\sum^{T-1}_{t=0}\left[D_{KL}\left(q_{\phi}(z_{t+1}|x_{t+1},z_{t},a_{t})\|p_{\phi}(z_{t+1}|z_{t},a_{t})\right)\right]\right]$ (12)
s.t. $\quad\mathop{\mathbb{E}}_{(x_{1:T},a_{1:T-1})\sim D}\left[\mathop{\mathbb{E}}_{z_{1:T}\sim q_{\phi}}\left[\sum^{T-1}_{t=0}-\log p_{\phi}(x_{t+1}|z_{t+1})\right]\right]\leq\xi_{\mathrm{score}}$ (13)
$\mathop{\mathbb{E}}_{(x_{1:T},a_{1:T-1},y_{1:T-1})\sim D}\left[\mathop{\mathbb{E}}_{z_{1:T}\sim q_{\phi}}\left[\sum^{T-2}_{t=0}-\log p_{\phi}(y_{t+1}|z_{t+1})\right]\right]\leq\xi_{\mathrm{adherence}}$ (14)
where in Eq. (13) we have regression with Gaussian distribution, and in Eq. (14) we use cross-binary entropy loss for classification. To solve the optimization problem, we incorporate the constraints into the objective function using Lagrange multipliers $\lambda$. We apply methods from (Chen et al., 2022) to adapt $\lambda$. This allows the model to balance the importance of the constraints relative to the divergence terms, which can help in avoiding common pitfalls in training such as suboptimal local minima and posterior collapse.
To avoid over-fitting, we incorporate dropout (Srivastava et al., 2014) and Mixup (Zhang et al., 2017). Subsequent research has extended the application of Mixup to latent variable models, specifically within the latent space (e.g., (Chen et al., 2020)). However, considering our need for data augmentation across all data dimensions, not limited to latent variables, we have selected to implement the original Mixup method in our experiments.
4.2.3 LSTM
Long short-term memory (LSTM) is a classical sequential model (Hochreiter and Schmidhuber, 1997). The primary objective of this study is to forecast $y_{t}$ utilizing historical data, formulated as $y_{t}=f(x_{1:t},y_{1:t-1},s)$. To align this approach with the SLAC model for score prediction, an additional term $x_{t+1}$ is also predicted,
$(x_{t+1},y_{t})=f(x_{1:t},y_{1:t-1},s)$ (15)
where $f$ is a function represented by LSTM. The loss consists of cross entropy for adherence classification and Normalized Mean Squared Error Loss (NMSE) for score prediction.
In our scenarios, SLAC stands out due to its inherent flexibility over traditional sequential models like LSTM. This flexibility is primarily observed in its predictive capabilities. SLAC can predict $y_{t}$ and use this prediction to influence the subsequent $x_{t+1}$. In contrast, LSTM only predicts a pair of $y_{t}$ and $x_{t+1}$ simultaneously, implying that we cannot use $y_{t}$ to alter $x_{t+1}$. Although it is possible to modify the LSTM model to predict a pair of $y_{t}$ and $x_{t}$, this approach encounters a similar issue for $y_{t}$: it cannot predict $y_{t}$ using the information from $x_{t}$.
5 Results
[…]
The change of nasal resistance (NR) and peak nasal inspiratory flow (PNIF) after nasal allergen provocation (NPT) was used to evaluate the severity of symptoms by combining the symptom score. The change of NR after NPT from the adherent group was higher than the non-adherent group $(30\,\%\text{ vs }14.8\,\%)$. The laboratory tests such as total IgE, sIgE of Derp and Derf, and SPT did not exhibit a significant difference between the two groups. For detailed characteristics of patients see Table 1.
The observed total non-adherence rate at the end of three years was 35.4% and the median of the SCIT duration was 18 months in the study. The rate of dropout in the third year(43.0%) was highest in comparison to the end of the first year $(26.5\,\%)$ and the second year $(30.0\,\%)$. The reason for the withdrawal from the patients included the concern of COVID-19, especially at the beginning of 2020 accounting for a $25\,\%$ portion of the non-adherent patients in the first year. The most influential reason for the withdrawal was unreached expectations for clinical improvement (43.9%). Medical issues including pregnancy status during the treatment period and other physical disorders were collected from patients leading to the withdrawal of SCIT $(4.5\,\%)$. The significantly improved symptoms contributed to the reason for dropout, especially after two years SCIT treatments $(18.2\,\%)$. The recorded cases from side effects accounted for $4.5\,\%$, comprising the local and systematic adverse reactions (see Table 2).
We have a total of 205 samples, which we have randomly divided into a test dataset comprising $20\,\%$, i.e., 41 samples. For our analysis, we employ a five-fold cross-validation approach. Additionally, we apply zero-mean and unit standard deviation (STD) normalization to the variables $x$ and $s$.
The Root Mean Square Error (RMSE) metric is used to evaluate the precision of our medical score predictions. Furthermore, to assess the adherence predictions, we utilize a comprehensive set of metrics including accuracy, precision, recall, and the F1 score, each offering a unique perspective on the performance of our predictive models. Most of the figures in this study are presented using boxplots.
The uncertainties for both models are calculated using five-fold cross-validation. In addition, as SLAC is a probabilistic model, we also perform 100 samples from the latent space to compute its uncertainty.
5.1 One-Step Prediction
Figure 3: RMSE of the prediction step by step. The red dashed line is the RMSE of random prediction with Uniform distribution. See Fig. 5 and 6 for more details.
Figure 4: Accuray of the prediction step by step. The red dashed line is the accuracy of random prediction with Uniform distribution. See Table 3 for more details.
Figure 5: RMSE of SLAC one-step prediction across various scores and time steps.
Figure 6: RMSE of LSTM one-step prediction across various scores and time steps.
metric model time step 1 time step 2 time step 3 time step 4 time step 5
accuracy LSTM $\mathbf{1.00\pm 0.00}$ $0.66\pm 0.03$ $\mathbf{0.80\pm 0.04}$ $\mathbf{0.84\pm 0.05}$ $\mathbf{0.74\pm 0.02}$
SLAC $1.00\pm 0.01$ $\mathbf{0.70\pm 0.06}$ $0.72\pm 0.04$ $0.71\pm 0.04$ $0.60\pm 0.06$
precision LSTM $\mathbf{1.00\pm 0.00}$ $0.72\pm 0.01$ $\mathbf{0.86\pm 0.06}$ $\mathbf{0.90\pm 0.05}$ $\mathbf{0.62\pm 0.03}$
SLAC $\mathbf{1.00\pm 0.00}$ $\mathbf{0.75\pm 0.03}$ $0.74\pm 0.03$ $0.71\pm 0.03$ $0.44\pm 0.05$
recall LSTM $\mathbf{1.00\pm 0.00}$ $0.86\pm 0.06$ $0.83\pm 0.05$ $0.82\pm 0.08$ $0.61\pm 0.03$
SLAC $1.00\pm 0.01$ $\mathbf{0.87\pm 0.06}$ $\mathbf{0.90\pm 0.03}$ $\mathbf{0.86\pm 0.04}$ $\mathbf{0.70\pm 0.10}$
F1 score LSTM $\mathbf{1.00\pm 0.00}$ $0.79\pm 0.03$ $\mathbf{0.84\pm 0.03}$ $\mathbf{0.85\pm 0.05}$ $\mathbf{0.62\pm 0.03}$
SLAC $1.00\pm 0.00$ $\mathbf{0.81\pm 0.04}$ $0.81\pm 0.02$ $0.78\pm 0.03$ $0.54\pm 0.06$
Table 3: Comparison of LSTM and SLAC over different time steps. The results are expressed as a mean ± standard deviation. The better results are highlighted in bold.
In this experiment, our focus is on predicting the immediate next step. Within the SLAC, the prediction of $y_{t}$ is based on the sequence $x_{1:t}$ and actions $a_{1:t-1}$. Additionally, we forecast the subsequent state $x_{t+1}$ using the sequence $x_{1:t}$ along with actions $a_{1:t}$. In contrast, for LSTM, the predictions for both $y_{t}$ and the next state $x_{t+1}$ are derived from $x_{1:t}$ and $y_{1:t-1}$.
As illustrated in Fig.3, SLAC surpasses LSTM in performance beginning at time step two. The figure indicates that with an increased amount of historical data (additional time steps), SLAC achieves greater RMSE. Both SLAC and LSTM demonstrate considerably better over random prediction methods. Further insights are provided in Fig. 5 and Fig. 6, which provides detailed representations of each feature. In the prediction of specific local symptoms score, SLAC performs an improvement in error after step two with all parameters compared to LSTM. The results from RMSE in nasal and ocular symptoms display relatively high values compared to those for lower respiratory tract symptoms. This can be attributed to the fact that the majority of patients in the cohort predominantly exhibited nasal and ocular symptoms, which presented a wide range of scores.
Fig. 4 demonstrates that from steps two to four, accuracy in adherence predictions improves with the inclusion of additional information. The first step exhibits a notable bias, as it only includes data from adherent patients, as detailed in Sec. 4.1.2. Nonetheless, both models adeptly manage this bias and achieve high-accuracy predictions. Prediction for the sixth step is not conducted due to the cessation of treatment by the hospital. In the fifth step, there is a decline in accuracy, likely due to the extended time interval of 12 months. Table 3 illustrates details of the classification for one-step prediction. Initially, both models exhibit perfect performance in Accuracy, Precision, and Recall at the first time step, but diverge in subsequent steps. In terms of Accuracy, LSTM generally outperforms SLAC, particularly evident at time steps three, four, and five. For Precision, LSTM again shows superior performance in the later time steps, except at time step two where SLAC marginally leads. However, in the Recall metric, SLAC surpasses LSTM from time step two onwards, indicating its strength in correctly identifying positive cases. The F1 score, which balances precision and recall, shows LSTM generally ahead, except at time step two where SLAC has a slight edge. This metric indicates LSTM’s balanced capability in both precision and recall, especially in the later time steps. Overall, while both models start equally strong, LSTM demonstrates greater consistency and effectiveness across most metrics and time steps. SLAC, while lagging slightly behind in accuracy and precision, shows its robustness in recall, especially in the middle to later time steps.
5.2 Rollouts
Figure 7: RMSE of the rollout prediction. The first time step in each subplot represents the beginning of the rollout time step.
Figure 8: Accuracy of the rollout prediction. The first time step in each subplot represents the beginning of the rollout time step.
In the rollout experiment, our focus extends to a longer-term prediction. The SLAC prediction of $y_{t:T-1}$ and $x_{t+1:T}$ are computed based on $x_{1:t}$ and $a_{1:t-1}$. Actions, $\{a_{i}:i\geq t\}$ are inferred from the model’s output, $y_{t}$. Moreover, for time steps greater than $t$, we employ a prior in the latent space, which eliminates the need for the input of $x_{t:T}$. In the LSTM model, the predictions for $y_{t:T-1}$ and $x_{t+1:T}$ are based on $x_{1:t}$ and $y_{1:t-1}$.
Fig. 7 and 8 illustrate the performance of our model in multi-step predictions. Similar to one-step predictions, the accuracy generally improves with the availability of more information, except in the case of the adherence prediction at the fifth step. The results demonstrate the model’s proficiency in making long-term predictions.
5.3 Model as a simulator
Given the initial condition of a patient, we can assess the outcomes of various interventions. Clinically, if the patient’s adherence to treatment significantly impacts the prognosis (and there is a possibility of non-adherence), it becomes imperative for the doctor to emphasize treatment compliance. Conversely, if adherence makes little difference, it suggests the therapeutic approach may be ineffective for this patient, allowing the doctors to emphasize adherence efforts.
To evaluate the impact of varying actions on SLAC’s performance, we analyze how different actions affect the resulting scores. In the absence of a ground truth with diverse actions for the same patient, our focus shifts to examining whether the states are responsive to changes in actions. Considering initial states $x_{1:3}$ and actions $a_{1:2}$, we do rollouts with $a_{3:5}$, alternating between one and zero. This controlled alteration reveals that the average predicted value of $x_{6}$ under these conditions is $-0.20$. This value is computed from the prediction outcomes for actions with ones minus those for actions with zeros. The result indicates that our model can be used as a simulator for doctors to see the impact of different treatments/therapies. Since the LSTM does not have similar functions (see Sec. 4.2.3), we only show the SLAC results.
5.4 Interpretability
Figure 9: Importances of the factors.
[…]
5.5 Architecture and computation
In this study, computational experiments were performed using an NVIDIA GeForce GTX 1080 Ti GPU, with the implementation done in PyTorch, version $2.1.0$.
The SLAC model’s architecture featured 32 hidden dimensions each for variables $z_{1}$ and $z_{2}$. Its encoder and decoder were symmetrically structured, each comprising five layers with 128 units. The primary activation function was LeakyReLU, set with a negative slope coefficient of 0.2. Both the encoder and decoder’s mean output layers were linear, while the STD layer utilized a Softplus activation. For binary classification tasks, a Sigmoid activation was used for output.
The LSTM architecture included a hidden dimension size of 128, with two LSTM layers. The output activation function for score prediction was linear, and like the SLAC model, a Sigmoid function was employed for binary classification outputs.
Both models shared the same optimization settings. They utilized the RAdam (Liu et al., 2019) optimizer with a learning rate of $0.001$. The batch size was set at 64, and a gradient clipping value of 0.8 was applied to ensure training stability. To prevent overfitting and enhance model generalization, a dropout rate of 0.05 was introduced. Additionally, both models incorporated Mixup as a data augmentation during training.
6 Discussion
The reported adherence rates of SCIT ranged from around $23\,\%$ to $90\,\%$, due to the non-uniform follow-up duration (2-4 years) (Passalacqua et al., 2013; Lee et al., 2019; Lemberg et al., 2017; Yang et al., 2018). Poor adherence in the three to five-year time span of AIT is an obstacle to reaching allergen tolerance and symptom remission. Recently proposed an ‘adherence and persistence in AIT (APAIT)’ checklist that assists researchers in assessing adherence or persistence to AIT in retrospective studies (Pfaar et al., 2023). The present study is the first research about the application of machine learning models in the adherence prediction of SCIT in AR patients. From our study, the accomplishment rate of the three-year treatment cycle was relatively low ($35.4\,\%$), while the rate of dropout after two years accounts for half ($42.8\,\%$) in the whole non-adherence cohort. Several researchers focus on these variables impacting the compliance of medical behaviors to enhance the intervention approach to reduce the withdrawal caused by disease-unrelated reasons. Even though the Covid-19 pandemic impacted the compliance of the majority of the patients in the three-year cycle, the first year since the pandemic’s outbreak appears to fundamentally built a barrier to patients, similar to the finding from Liu et al. (2021) that $11\,\%$ dropouts in the two years SCIT was observed caused by Covid-19. We excluded the patients who dropped out in the dose buildup phase within four months to minimize the dose-origin impact. Due to the uncovered cost from the public health care system and commercial insurance, financial burden accounts for a non-negligible factor in influencing the decision-making of patients. A similar finding from Lourenço et al. (2020). indicated that economic reasons contributed to the most frequent cause of SCIT cessation.
In comparison to the logistic regression analyses in the clinical field to identify independent predictors, the model from our study performed convincing interpretability. Previous research primarily concentrated on non-sequential prediction methods for adherence (Mousavi et al., 2022; Wang et al., 2020; Warren et al., 2022; Ruff et al., 2019). This approach presents a significant limitation in treatment processes, particularly for immunotherapy that often spans extended periods, such as three years. These non-sequential methods tend to predict only the overall outcome, overlooking the intricacies of intermediate time steps. To facilitate earlier intervention, a sequential model capable of making predictions at any given time step would be markedly more beneficial. While some subsequent studies have introduced sequential models (Hsu et al., 2022; Singh et al., 2022; Schleicher et al., 2023), their scope was restricted to predicting adherence alone. Our study enhances this approach by incorporating a state-action model, which can predict both adherence and score/state. This advancement allows for more precise and detailed analysis of patient cases by medical professionals.
Our study demonstrates notable findings in the domain of patient adherence prediction in subcutaneous immunotherapy. The comparison between the SLAC model and LSTM model reveals the distinct strengths and limitations of each approach. Notably, the SLAC exhibits greater flexibility, and it outperforms the LSTM in score prediction. This advantage likely stems from its ability to efficiently learn and generalize in complex environments, a trait that is crucial in medical data analysis. Conversely, the LSTM model shows better performance in predicting adherence, indicating its potential utility in scenarios where accurate forecasting of patient compliance is critical. Both models demonstrate the capability to handle longer sequences, extending beyond one-step prediction. This ability is crucial in medical settings where long-term patient monitoring and prediction are essential for effective treatment planning.
Overall, the study underscores the importance of selecting the appropriate model based on the specific requirements of the task, whether it be flexibility, precision in score prediction, or adherence prediction. The findings contribute to the growing field of machine learning applications in healthcare, particularly in enhancing patient-centered treatment strategies through accurate and personalized predictions. Future research could focus on evaluating the SLAC model’s performance in simulating various actions, further enriching its applicability in clinical settings.
7 Conclusion
In summary, our study showcases the flexibility and score prediction superiority of the Sequential Latent Actor-Critic (SLAC) model, while highlighting the superior adherence prediction capabilities of the Long Short-Term Memory (LSTM) model. Both models demonstrate proficiency in managing extended sequence predictions, extending beyond single-step forecasts. This study emphasizes the utility and significance of selecting the appropriate model based on specific task requirements, be it for precision in score prediction or adherence forecasting. These findings make a substantial contribution to the application of machine learning in healthcare, particularly in improving patient-centered treatment strategies through accurate and tailored predictions.
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Data Availability Statement
The dataset for this study can be found in the Subcutaneous-Immunotherapy-Dataset repository.